OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1).

OpenStudy (anonymous):

solve for y first :)

OpenStudy (anonymous):

The equation of the line is : \[\Large y=-3x+7\] So the slope is -3 The slope m of a perpendicular line to this line verify : \[\Large m\times(-3)=-1\] So : \[\Large m=\frac13\] So the equation is : \[\Large y=mx+b=-\frac13x+b\] For finding b, we have the point (6.-1) is from the line so : \[\Large -1=\frac13\times 6+b\] So : \[\Large b=-1-2=-3\] So the equation of line : \[\Large y=-\frac13x-3\] Or : \[\Large 3y+x+9=0\]